a group of 80 people attend a ball game. there were four times as many children as adults in the group. write a system of equations that you could use to set up this problem, where a is the number of adults and c is the number of children in the group. solve the system of equations for c, the number of childern in the group
Answer by shahid(44)

80 people are there so first equation is
a+c=80....(1)
second equ formed by condition is
c=4a ....(2) subtacting both equations
a=80-4a
a+4a=80
5a=80
a=80/5
a=16 so c=4(16)=64 there were 64 children in the group